OpenStudy (anonymous):

solve for x: (2x-5)^{2}=11

OpenStudy (rosey):

i nakey :)

OpenStudy (anonymous):

U can write this eq as 2x -5 = (11)^1/2 And now solve for x.. Don't worry abt roots..... U should Know the right method :)

jhonyy9 (jhonyy9):

4x^2 -20x +25 =11 4x^2 -20x +14=0 /2 2x^2 -10x +7=0 10+/- (radical 44) 2(5+/-radical 11) 5+/- radical 11 x_1,_2 = ---------------- = ---------------- = -------------- 4 4 2

OpenStudy (anonymous):

what

jhonyy9 (jhonyy9):

this is how do you need to do

jhonyy9 (jhonyy9):

- so but you can do it how have arote to you firstly Ishaan94

OpenStudy (anonymous):

squeaky 22 i m trying to upload a file which is ur answer sorry it is not uploading

jhonyy9 (jhonyy9):

- than you will get those results too

OpenStudy (anonymous):

squeeaky What problem are u facing?

OpenStudy (anonymous):

my results were different

jhonyy9 (jhonyy9):

write here how you have got those results

jhonyy9 (jhonyy9):

Nancy - pardon but i think that you need to divid +/- radical 11 with2 too not just -only - 5 - check it please newly

OpenStudy (anonymous):

\[(2x-5)^2=11\]\[2x-5=\pm \sqrt{11}\]\[x= \frac{5\pm \sqrt{11}}{2}\]